OpenStudy (superman2):

Find the length of arc AC Express your answer in terms of pi AB = 12 AND dEGREE Is 60

OpenStudy (superman2):



OpenStudy (iroman12):

where is arc C?

OpenStudy (superman2):

whoops sorry

OpenStudy (iroman12):

I mean point C in the circle?

OpenStudy (superman2):



OpenStudy (superman2):

sorry

OpenStudy (iroman12):

its okay

OpenStudy (superman2):

Formula for the length of an arc is : Pi times diameter i might be getting my fourmlas mix up

OpenStudy (iroman12):

so since the angle given is 60 degrees, we assume the line that you draw will be equal to 180 degrees. To find the angle of arc AC, we subtract 180 - 60 which is 120 degrees. We will be using this later. In a circle, it holds true that in any points of the circle represent the radius, so to find the radius of the circle given, \[a=r \theta \] where \[r=a/\theta \] where the value of theta is in radians not degrees \[r=12/{\Pi/3} \] which is equal to 11.5 feet. We will be using this radius to find the measurement of arc AC. \[a=r \theta \] \[a=(11.5)(2\Pi/3)\] which is equal to 24.1 feet. Therefore, arc AC is equal to approximately 24 feet.

OpenStudy (wolf1728):

I think they want the answer in terms of PI radius = 6 circumference (entire circle) = 2*PI*6 circumference for 60 degrees = 1/6 circumference 60 degrees = 2*PI*6 * (1/6) circumference 60 degrees = 2 * PI

OpenStudy (superman2):

So 2 pi the answer?

OpenStudy (wolf1728):

Yes, 2 * PI is the answer.